From a bike shop, Rama pedals due north while Sam rides due west. When Sam is 4 mi from the shop, the distance between her and Rama is two miles more than Rama’s distance from the bike shop. Find the distance between Rama and Sam.
Answer by Boreal(15235)

x is the distance from R to the bike shop and one leg of a right triangle (draw this)
Sama is 4 miles away (another leg).
The hypotenuse of the right triangle is the distance between them of x+2 miles
so 4^2+x^2=(x+2)^2
or x^2+4x+4=x^2+16
4x=12
x=3 miles
the distance between them is x+2 or 5 miles (3-4-5 right triangle)